### v2

# Multi-modal Stance Detection: New Datasets and Model

Abstract
Stance detection is a challenging task that aims to identify public opinion from social media platforms with respect to specific targets. Previous work on stance detection largely focused on pure texts. In this paper, we study multi-modal stance detection for tweets consisting of texts and images, which are prevalent in today’s fast-growing social media platforms where people often post multi-modal messages. To this end, we create five new multi-modal stance detection datasets of different domains based on Twitter, in which each example consists of a text and an image. In addition, we propose a simple yet effective Targeted Multi-modal Prompt Tuning framework (TMPT), where target information is leveraged to learn multi-modal stance features from textual and visual modalities. Experimental results on our three benchmark datasets show that the proposed TMPT achieves state-of-the-art performance in multi-modal stance detection.
Multi-modal Stance Detection: New Datasets and Model
Bin Liang1,2,4∗, Ang Li1,2thanks:  The first two authors contribute equally to this work., Jingqian Zhao1,2, Lin Gui5, Min Yang6, Yue Yu3, Kam-Fai Wong4, and Ruifeng Xu1,2,3thanks:  Corresponding Author 1 Harbin Institute of Technology, Shenzhen, China 2 Guangdong Provincial Key Laboratory of Novel Security Intelligence Technologies 3 Peng Cheng Laboratory, Shenzhen, China 4 The Chinese University of Hong Kong, Hong Kong, China  5 King’s College London, UK 6 SIAT, Chinese Academy of Sciences, Shenzhen, China bin.liang@cuhk.edu.hk,  angli@stu.hit.edu.cn,  xuruifeng@hit.edu.cn
1 Introduction
Stance detection is an important task for learning public opinion from social media platforms, which aims to determine people’s opinionated standpoint or attitude (e.g., Favor, Against, or Neutral, etc.) expressed in the content towards a specific target, topic, or proposition Somasundaran and Wiebe (2010); Augenstein et al. (2016). Existing conventional machine learning-based methods Hasan and Ng (2013); Mohammad et al. (2016); Ebrahimi et al. (2016) and deep learning-based methods Augenstein et al. (2016); Sun et al. (2018); Zhang et al. (2020); Chen et al. (2021); Allaway et al. (2021); Liang et al. (2022a) have made promising progress in different types of stance detection tasks for pure texts.
However, more and more present-day social media platforms like Twitter allow people to post multi-modal messages, which encourages people to express their stances and opinions through multi-modal content, posting texts with images for example. That is, detecting stance from the pure text modality may not accurately identify the user’s real view of a target. For example, Figure 1 shows a post composed of a text and an image. The stance expression towards “Donald Trump” and “Joe Biden” in this example can not be accurately identified based on text information unless combined with the information of the visual modality. Therefore, how to detect users’ stances on a topic from multi-modal posts might help better identify public opinion in social media.
Figure 1: An example of a user expressing an “Against” stance on “Donald Trump” and a “Favor” stance on “Joe Biden” using multi-modal information.
For multi-modal stance identification, Weinzierl and Harabagiu (2023) discussed the multi-modal stance on frames of communication. However, this work ignored the consideration of stance towards a target, which is an important factor in stance detection. Therefore, aiming to push forward the research of multi-modal stance detection, we create five new datasets, in which each example consists of a specific target, a text, and an image. These datasets contain a total of 17,544 annotated data entries across 5 domains and 12 targets, including hot topics, politicians, and debates.
To deal with multi-modal stance detection, we propose a simple yet effective Targeted Multi-modal Prompt Tuning framework (TMPT), where the targeted prompt tuning is employed to adapt pre-trained models for learning stance features from different modalities. Specifically, to leverage the target information in stance detection, we first devise targeted prompts for both textual and visual modalities. Then, the targeted prompts are fed to the pre-trained language model and pre-trained visual model to learn stance features for the target from different modalities. Further, a simple vector concatenation is used to fuse the features from different modalities for multi-modal stance detection.
The main contributions of our work are summarized as follows:
1) We manually annotate five new multi-modal stance detection datasets based on Twitter data from different domains. The release of the datasets would push forward the research in this field.
2) A simple yet effective targeted multi-modal prompting tuning framework is proposed to deal with multi-modal stance detection, where the target information is used to prompt the pre-trained models for learning multi-modal stance features.
3) A series of experiments on our datasets show that the proposed method significantly outperforms the baseline models [^1] To facilitate future research, all datasets and code are publicly available at https://github.com/Leon-Francis/Multi-Modal-Stance-Detection.
Dataset Target # Samples and Proportion of Labels
Mtse Favor % Against % Neutral % - - Total
Donald Trump (DT) 231 14.03 1297 78.75 119 7.23 - - 1647
Joe Biden (JB) 602 47.78 524 41.59 134 10.63 - - 1260
Mccq Favor % Against % Neutral % - - Total
Chloroquine (CQ) 455 33.58 503 37.12 397 29.30 - - 1355
Mwtwt Support % Refute % Comment % Unrelated % Total
CVS_AET 426 24.38 65 3.72 866 49.57 390 22.32 1747
CI_ESRX 321 35.71 91 10.12 298 33.15 189 21.02 899
ANTM_CI 59 5.01 238 20.22 306 26.00 574 48.77 1177
AET_HUM 94 9.82 287 29.99 267 27.90 309 32.29 957
DIS_FOXA 417 13.97 103 3.45 1504 50.37 962 32.22 2986
Mruc Support % Oppose % Neutral % - - Total
Russia (RUS) 16 1.4 763 68.74 331 29.82 - - 1110
Ukraine (UKR) 742 68.64 39 3.61 300 27.75 - - 1081
Mtwq Support % Oppose % Neutral % - - Total
Mainland of China (MOC) 339 24.27 834 59.70 224 16.03 - - 1397
Taiwan of China (TOC) 1595 82.73 74 3.84 259 13.43 - - 1928
Table 1: Label distribution of the five multi-modal stance detection datasets.
2 Related Work
Textual Stance Detection
Various methods based on conventional machine learning Hasan and Ng (2014); Mohammad et al. (2016) and deep learning Sun et al. (2018); Zheng et al. (2022); Li and Caragea (2023); Li et al. (2023) have been proposed to deal with the stance detection regarding a specific target. For performing stance detection in real-world scenarios, many existing methods focus on the task of zero-shot stance detectionLiu et al. (2021b); Liang et al. (2022b); Wen and Hauptmann (2023); Zhao et al. (2023).
Prompt Tuning
Prompting Liu et al. (2021a) was originally aimed to design language instructions for pre-trained language models (PLMs) to transfer learning in downstream tasks Shin et al. (2020); Jiang et al. (2020). Recent works begin to treat prompts as continuous vectors and optimize them during fine-tuning, called Prompt Tuning Li and Liang (2021); Lester et al. (2021); Liu et al. (2021c). Besides, Radford et al. (2021); Zhou et al. (2022); Ju et al. (2022); Jia et al. (2022) introduce prompt into vision-language models to leverage the ability of prompt tuning in multi-modal tasks.
Image Person Events Words Memes Mixed
Mtse 43.7 32.7 8.2 24.0 25.0 10.1
Mccq 44.1 12.2 22.9 26.7 16.3 21.9
Mwtwt 46.2 26.4 16.8 38.4 5.9 12.5
Mruc 54.8 32.1 20.7 20.4 18.5 8.3
Mtwq 41.0 36.8 42.9 8.1 2.4 9.8
Table 2: The statistics of whether the image conveys stance information (%Image) and the type of each image (%Person, %Events, %Words, %Memes, %Mixed).
3 Multi-modal Stance Detection Datasets
Based on three open-source textual stance detection datasets: Twitter Stance Election 2020 Kawintiranon and Singh (2021), COVID-CQ Mutlu et al. (2020), and Will-They-Won’t-They Conforti et al. (2020), and two hot topics in recent years: Russo-Ukrainian Conflict [^2] https://en.wikipedia.org/wiki/Russo-Ukrainian_War and Taiwan Question [^3] https://en.wikipedia.org/wiki/Cross-Strait_relations. We create five multi-modal stance detection datasets of different domains to provide available data for this task: Multi-modal Twitter Stance Election 2020 (Mtse), Multi-modal COVID-CQ (Mccq), Multi-modal Will-They-Won’t-They (Mwtwt), Multi-modal Russo-Ukrainian Conflict (Mruc) and Multi-modal Taiwan Question (Mtwq).
3.1 Data Collection
We use TweetScraper [^4] https://github.com/jonbakerfish/TweetScraper to retrieve tweets with corresponding keywords of five datasets (shown in Appendix A), keeping the posts containing text in English and at least one image or video/GIF. For videos/GIFs, we retain their first frame. For posts with multiple images, we combine the text with each image to form multiple samples. Finally, we obtain 130000, 90000, 60000, 100000, and 80000 candidate examples of five datasets, respectively.
3.2 Data Annotation
To ensure consistency with previous stance detection work, we follow the guidelines of Twitter Stance Election 2020 Kawintiranon and Singh (2021), COVID-CQ Mutlu et al. (2020), and Will-They-Won’t-They Conforti et al. (2020) to annotate the multi-modal stance of Mtse, Mccq and Mwtwt. For Mruc and Mtwq, the annotation guidelines are shown in Appendix B. To facilitate the analysis of how visual modality affects stance detection in our constructed datasets, we also instruct annotators to indicate whether the image conveyed stance information. We invite eight experienced researchers [^5] We recruit experienced researchers who have worked on multi-modal learning over 3 years. to label the stance for each example. Each sample will be annotated by three different annotators, and the gold label is obtained by majority vote. For the disagreed results among the three annotators, we invited three additional annotators to annotate and then performed a majority vote to obtain the gold label [^6] During the annotation process, only 2.54% of the data need to be allocated to additional annotators..
3.3 Quality Assessment
We use Cohen’s Kappa Statistic to evaluate the inter-annotator agreement Cohen (1960). The average Cohen’s Kappa between our annotator pairs for Mtse is 0.703, for Mccq is 0.689, for Mwtwt is 0.729, for Mruc is 0.752, and for Mtwq is 0.691. This demonstrates that the Kappa scores of all datasets are substantial. In addition, the average Cohen’s Kappa reported in the related textual stance detection dataset Will-They-Won’t-They Conforti et al. (2020) is 0.67, which also indicates the high quality of our new datasets from another angle.
Figure 2: The overall architecture of our proposed TMPT. Textual Prompting is devised for adapting the large pre-trained language model. Visual Prompt Tuning is devised for adapting the large pre-trained vision model.
3.4 Data Analysis
We finally got 17,544 well-annotated multi-modal samples across 5 domains and 12 targets. Each example consists of a text in English and an associated image. The statistics of datasets are reported in Table 1. Note that differences in label distribution between targets are common, which is also observed in other textual stance detection datasets Mohammad et al. (2016); Mutlu et al. (2020); Conforti et al. (2020). Through analyzing the indication of whether the image conveys stance information introduced in Section 3.2, and analyzing the type of each image, we obtained statistics shown in Table 2. It indicates that in our five multi-modal stance detection datasets, nearly half of the samples will incorporate stance information due to the combination of visual modality, therefore facilitating stance detection. Further, the proportion of image types varies greatly among different datasets. This complicates image comprehension for stance detection, which is also the main challenge of multi-modal stance detection.
4 Methodology
This section details our proposed targeted multi-modal prompt tuning framework (TMPT). Given a text $S$ and an image $I$, the goal of multi-modal stance detection is to identify the stance label $y$ for the specific target $t$. Therefore, to leverage the target information in this task, we design targeted multi-modal prompt tuning for both textual and visual modalities, which are utilized to prompt the pre-trained models for learning multi-modal stance features. The architecture of our TMPT is shown in Figure 2, containing four main components: 1) Textual Prompt Tuning, which encodes the input of textual modality based on the textual targeted prompt; 2) Visual Prompt Tuning, which encodes the input of visual modality based on the visual targeted prompt; 3) Multi-modal Fusion, which fuses the representations from textual and visual modalities to capture the stance features. 4) Multi-modal Stance Detection, which derives the stance label for an input example according to the multi-modal stance features.
Dataset Target Textual Targeted Prompt
Mtse DT The stance on Donald Trump is:
Mccq CQ The stance on the use of Chloroquine and Hydroxychloroquine for the treatment or prevention from the coronavirus or COVID 19 is:
Mwtwt CVS_AET The stance on merger and acquisition between CVS Health and Aetna is:
Mruc RUS The stance on Russia is:
Mtwq MOC The stance on Mainland of China is:
Table 3: The example of textual targeted prompts.
4.1 Textual Prompt Tuning
Textual Prompt Construction
Inspired by the textual prompt tuning Liu et al. (2023), considering the characteristics of stance detection, we devise a textual targeted prompt for each text to adapt the pre-trained language model to the stance detection task. As the example shown in Figure 2, take the target “Donald Trump” as an example, the textual targeted prompt is defined as:
$\begin{split}\mathcal{P}^{T}=\mathrm{The~{}stance~{}on~{}Donald~{}Trump~{}is:}\end{split}$ (1)
That is, the textual targeted prompts are designed according to the targets and the purpose of stance detection. The examples of textual targeted prompts of other targets are shown in Table 3. In Section C.1, we introduced different textual prompt settings.
Textual Encoder
Based on the textual targeted prompts, we construct the input of each text on target $t$. Given a text $S$ consists of a sequence of words $S=\{w_{i}\}_{i=1}^{n}$, $n$ is the length of $S$. The input of the textual modality is represented as:
[…]
Then, we adopt the pre-trained uncased BERT-base model Devlin et al. (2019) to map each textual token into a $d^{T}$-dimensional embedding:
$[{\bm{x}}^{T}_{\texttt{[CLS]}},{\bm{x}}^{T}_{1},\cdots,{\bm{x}}^{T}_{m+n+2}]={\rm BERT}(\mathcal{I}^{T})$ (3)
Where ${\bm{x}}^{T}_{\texttt{[CLS]}}$ represents the embedding of the [CLS] token, $m$ is the length of the textual targeted prompt. In this way, the pre-trained language model can encode the input text according to the target, and obtain the feature representation that contains the target-specific stance information.
4.2 Visual Prompt Tuning
Visual Embedding Layer
Following Dosovitskiy et al. (2021), we first divide the image $I$ into $r$ fixed-sized patches $I=\{{\bm{p}}_{i}\in\mathbb{R}^{l^{2}\times c}\}_{i=1}^{r}$, where $(l,l)$ is the resolution of each patch, $c$ is the number of channels. Then, following Jia et al. (2022), we embed each patch into a $d^{V}$-dimensional vector with positional encoding ${\bm{E}}_{pos}$ by the embedding projection ${\bm{E}}$:
${\bm{v}}_{j}={\bm{p}}_{j}{\bm{E}}\quad\quad{\bm{v}}_{j}\in\mathbb{R}^{d},j\in\{1,r\}$ (4)
${\bm{V}}^{0}=[\{{\bm{v}}_{j}\}_{j=1}^{r}]+{\bm{E}}_{pos}$ (5)
Where ${\bm{V}}^{0}$ is the embedding for the input image.
Visual Prompt Construction
Inspired by the visual prompt tuning proposed by Jia et al. (2022), we devise visual targeted prompt, aiming to instruct the pre-trained vision model to learn the features according to the specific target. Specifically, we introduce continuous embedding (visual prompt tokens), as the visual prompt for target $t$. Each prompt consists of $\lambda$ learnable embedding, which can be formulated as follows:
$\mathcal{P}^{V}=\{{\bm{e}}_{i}\in\mathbb{R}^{d^{V}}|1\leq i\leq\lambda\}$ (6)
Here, each different target corresponds to a different set of visual prompt embedding. In Section C.2, we introduced different initialization methods of visual prompt embedding.
Visual Encoder
Based on the visual embedding ${\bm{V}}^{0}$ and the visual targeted prompts $\mathcal{P}^{V}$, we use the pre-trained Vision Transformer model ViT Dosovitskiy et al. (2021) with $N$ layers to encode the input image for learning visual stance features on target $t$. Here, the targeted prompts are inserted into the first Transformer layer $L^{1}$. Therefore, the first layer targeted prompted ViT is defined as:
$[{\bm{x}}^{V}_{\texttt{[CLS]}_{1}},{\bm{Z}}_{1},{\bm{V}}_{1}]=L^{1}([{\bm{x}}^{V}_{\texttt{[CLS]}_{0}},\mathcal{P}^{V},{\bm{V}}_{0}])$ (7)
For layer $k\in\{2,N\}$, the targeted prompted ViT is defined as:
$[{\bm{x}}^{V}_{\texttt{[CLS]}_{k}},{\bm{Z}}_{k},{\bm{V}}_{k}]=L^{k}([{\bm{x}}^{V}_{\texttt{[CLS]}_{k-1}},{\bm{Z}}_{k-1},{\bm{V}}_{k-1}])$ (8)
Here, we use ${\bm{x}}^{V}_{\texttt{[CLS]}}$ to represent the embedding of [CLS] token learned by the final Transformer layer $N$, i.e., ${\bm{x}}^{V}_{\texttt{[CLS]}}={\bm{x}}^{V}_{\texttt{[CLS]}_{N}}$. By using the visual prompts corresponding to each target, the pre-trained visual model can encode the input image according to the target, and obtain the feature representation that contains the target-specific stance information.
4.3 Multi-modal Fusion
Based on the embeddings of [CLS] tokens, we first use two dense layers with Leaky ReLU Maas et al. (2013) to derive the hidden representations of textual and visual modalities:
${\bm{h}}^{T}=f({\bm{W}}^{T}{\bm{x}}^{T}_{\texttt{[CLS]}}+{\bm{b}}^{T})$ (9)
${\bm{h}}^{V}=f({\bm{W}}^{V}{\bm{x}}^{V}_{\texttt{[CLS]}}+{\bm{b}}^{V})$ (10)
Where ${\bm{h}}^{T}\in\mathbb{R}^{d^{h}}$ and ${\bm{h}}^{V}\in\mathbb{R}^{d^{h}}$ are the hidden representation of textual and visual modalities respectively. $d^{h}$ is the dimensionality of the hidden representation. $f(\cdot)$ represents the Leaky ReLU activation function. ${\bm{W}}^{T}\in\mathbb{R}^{d^{h}\times d^{T}}$ and ${\bm{W}}^{V}\in\mathbb{R}^{d^{h}\times d^{V}}$ are weight matrices. ${\bm{b}}^{T}\in\mathbb{R}^{d^{h}}$ and ${\bm{b}}^{V}\in\mathbb{R}^{d^{h}}$ are biases.
In our TMPT, vector concatenation is used to fuse the feature vectors from different modalities:
${\bm{h}}={\bm{h}}^{T}\oplus{\bm{h}}^{V}$ (11)
Where ${\bm{h}}\in\mathbb{R}^{2d^{h}}$ is the final multi-modal stance representation. $\oplus$ represents the vector concatenation operation.
4.4 Multi-modal Stance Detection
Then, the final multi-modal stance representation is fed into a fully-connected layer with a softmax function to capture a probability distribution $\hat{\bm{y}}\in\mathbb{R}^{d^{p}}$ in the stance decision space:
$\hat{\bm{y}}={\rm{softmax}}({\bm{W}}^{o}{\bm{h}}+{\bm{b}}^{o})$ (12)
Where ${d^{p}}$ is the dimensionality of stance labels. ${\bm{W}}^{o}\in\mathbb{R}^{d^{p}\times 2d^{h}}$ and ${\bm{b}}^{o}\in\mathbb{R}^{d^{p}}$ are weight matrix and bias respectively.
Modality Method MTSE MCCQ MWTWT MRUC MTWQ
DT JB CQ CA CE AC AH DF RUS UKR MOC TOC
Textual BERT 50.97 65.29 66.57 69.56 65.03 69.74 67.99 79.21 48.44 64.32 63.44 54.32
RoBERTa 51.67 65.03 67.93 69.31 66.32 71.62 66.01 77.41 48.74 67.64 65.96 54.40
KEBERT 56.81 68.26 66.84 71.67 67.56 69.29 69.74 80.57 49.11 65.20 72.22 61.67
[…]
GPT4 68.74 66.39 65.84 63.14 65.12 69.93 71.62 52.69 47.04 50.55 68.47 45.05
Visual ResNet 37.89 38.59 47.16 39.89 42.20 43.52 37.05 50.34 39.49 39.28 44.92 38.60
ViT 40.48 40.42 46.64 46.63 50.00 40.16 46.32 50.86 38.44 41.28 46.93 36.28
SwinT 39.89 40.43 48.80 46.30 46.99 41.02 47.39 51.32 39.56 41.75 48.28 39.08
Multi- modal BERT+ViT 52.88 66.29 69.32 70.30 66.17 70.11 70.02 80.81 47.93 53.71 72.25 56.81
ViLT 35.32 48.24 47.85 62.70 56.44 58.06 60.22 73.66 45.16 50.59 56.12 59.51
CLIP 53.22 65.83 63.65 70.93 67.17 67.43 70.86 79.06 52.20 64.69 69.68 59.77
[…]
GPT4-Vision 70.46 72.82 61.63 44.59 47.07 57.47 57.90 37.61 44.27 55.01 68.93 51.61
[…]
TMPT+CoT 74.36 76.57 75.36 75.39 69.78 74.39 73.83 81.10 63.81 71.67 77.91 66.74
Table 4: Experimental results (%) of in-target multi-modal stance detection. The dark background results are for our TMPT. Best scores of each group are in bold. Second best scores are with underlines. Results with $\star$ denote the significance tests of our TMPT over the baseline models at $p\rm{-}value<0.05$.
5 Experimental Setup
To advance and facilitate research in the field of multi-modal stance detection, based on our new datasets, we conduct experiments on In-target Multi-modal Stance Detection, training and testing on the same target, and Zero-shot Multi-modal Stance Detection, performing stance detection on unseen targets based on the known targets.
5.1 Data Partition
For in-target multi-modal stance detection, each dataset is divided into training, development, and testing sets with a ratio of 70%:10%:20%. For zero-shot multi-modal stance detection, the training and development set is built on known target(s), and the testing set is built on unknown target(s). Since there is only one target in the Mccq dataset, we only use it for in-target multi-modal stance detection. In the Mwtwt dataset, following Conforti et al. (2020), we select four targets (CVS_AET, CI_ESRX, ANTM_CI, and AET_HUM) to perform zero-shot scenario since DIS_FOXA is not in the same domain as them.
5.2 Comparison Models
Pure textual modality baselines: 1) BERT Devlin et al. (2019), the uncased BERT-base; 2) RoBERTa Liu et al. (2019a), the RoBERTa-base; 3) KEBERT Kawintiranon and Singh (2022), a BERTweet-base model with specific knowledge of Twitter political posts. 4) LLaMA2 Touvron et al. (2023), the LLaMA2-70b-chat; 5) GPT4  [^7] https://openai.com/research/gpt-4. Pure visual modality baselines: 1) ResNet He et al. (2016), the ResNet-50 v1.5; 2) ViT Dosovitskiy et al. (2021), the vit-base-patch16-224; 3) Swin Transformer (SwinT) Liu et al. (2021d), the swinv2-base-patch4-window12-192-22k. Multi-modal baselines: 1) ViLT Kim et al. (2021), the vilt-b32-mlm; 2) CLIP Radford et al. (2021), the clip-vit-base-patch32; 3) BERT+ViT, utilizing BERT as the textual encoder and ViT as the visual encoder, and concatenating the [CLS] vectors of textual and visual modalities for stance detection. 4) Qwen-VL Bai et al. (2023), the Qwen-VL-Chat-7b. 5) GPT4-Vision  [^8] https://openai.com/research/gpt-4v-system-card.
5.3 Experimental Settings
To leverage the powerful capabilities of large language models, following Gatto et al. (2023), we also established a variant towards our TMPT: Input text, image, and chain-of-thought generated by GPT-4-Vision (TMPT+CoT). We utilize the pre-trained uncased BERT-base Devlin et al. (2019) to embed each word as a 768-dimensional embedding and employ the pre-trained ViT-base Dosovitskiy et al. (2021) to embed each image patch as a 768-dimensional embedding, i.e., $d^{T}=d^{V}=768$. The resolution of the visual patch is set to $(16,16)$. The length of visual prompt tokens is set to $c=7$. The dimensionality of hidden vectors is set to $d_{h}=768$. We use Macro F1-score to measure the model performance. The experimental results of our models are averaged over 5 runs to ensure the final reported results are statistically stable. For detailed settings of the experiments, please refer to Appendix C.
Modality Method MTSE MWTWT MRUC MTWQ
DT JB CA CE AC AH RUS UKR MOC TOC
Textual BERT 26.60 32.21 59.22 59.22 64.86 57.46 22.31 18.88 27.31 15.80
RoBERTa 27.23 30.57 61.67 63.93 66.28 60.93 25.10 20.53 29.32 11.46
KEBERT 26.17 31.81 59.70 62.56 63.92 55.53 24.28 23.39 28.91 20.84
[…]
GPT4 70.78 68.83 57.19 60.56 65.63 69.01 45.25 49.84 65.78 53.04
Visual ResNet 25.52 29.70 23.01 24.11 25.21 25.27 23.46 27.06 27.40 25.85
ViT 28.63 29.70 24.59 28.18 34.06 33.40 19.63 21.34 29.11 24.96
SwinT 28.54 30.85 28.53 28.50 35.87 34.33 25.13 24.39 28.30 21.19
Multi-modal BERT+ViT 26.70 31.57 59.21 59.30 65.04 59.28 24.69 22.83 26.74 20.68
ViLT 28.08 29.74 38.33 46.00 55.01 48.55 21.38 23.84 23.63 19.25
CLIP 28.21 28.99 61.08 55.67 63.80 60.06 26.29 26.39 27.15 16.33
[…]
GPT4-Vision 72.68 71.28 42.23 45.92 54.59 53.19 43.57 47.51 66.70 52.33
[…]
TMPT+CoT 45.58 56.16 66.75 68.69 67.28 63.78 52.24 57.40 54.17 48.90
Table 5: Experimental results of zero-shot multi-modal stance detection. The dark background results are for our TMPT. Best scores of each group are in bold. Results with $\star$ denote the significance tests of our TMPT over the baseline models at $p\rm{-}value<0.05$.
6 Experimental Results
6.1 In-target Multi-modal Stance Detection
The results of in-target multi-modal stance detection are shown in Table 4. It can be seen that our proposed TMPT outperforms existing baselines on all datasets, which verifies the effectiveness of our method. Further, TMPT+CoT performs overall better than TMPT, which demonstrates that leveraging the knowledge from large language models can improve the comprehension of textual modality, thereby achieving better performance. For uni-modal baselines, the performance of visual modality methods is unsatisfactory, while the performance of textual modality methods is much better. This indicates that the stance expression primarily resides in the textual modality. Further, models with visual modality perform better than single-text modality ones, which proves the importance of visual modality in multi-modal stance detection. This also implies how making better use of visual information is key to improving multi-modal stance detection performance. In addition, KEBERT, which integrates specific knowledge of Twitter political posts into BERTweet, achieves promising performance compared to other textual models, which indicates that exploring external target-related knowledge might improve the performance of multi-modal stance detection. Compared with BERT+ViT, ViLT, and CLIP, our TMPT achieves better performance, denoting that exploiting targeted prompt tuning can preferably leverage the target-specific multi-modal stance information, thus improving stance detection performance.
6.2 Zero-shot Multi-modal Stance Detection
The results of zero-shot multi-modal stance detection are reported in Table 5. It can be seen that the large language models achieve superior performance due to the need for detecting stances on unseen targets. This may be attributed to the powerful zero-shot learning ability of large language models For our TMPT, which does not use chain-of-thoughts from large language models, still achieves better performance than large language models in some targets, while also outperforming all non-large language model baselines. This shows the promise of our TMPT in multi-modal stance detection. Further, TMPT+CoT performs overall better than TMPT, which indicates that obtaining powerful text and visual comprehension abilities from large models may be key to improving the performance of detecting stance for unseen targets.
Method MTSE MCCQ MWTWT MRUC MTWQ
TMPT 66.66 71.36 79.51 60.01 72.33
[…]
w/o $\mathcal{P}^{V}$ 64.86 69.66 78.06 59.15 71.73
Table 6: Experimental results of ablation study. The reported results are the average results of all targets in a dataset on in-target multi-modal stance detection.
6.3 Ablation Study
To analyze the impact of the targeted prompt tuning in our proposed TMPT, we conduct an ablation study and report the results in Table 6. Note that the removal of textual prompting (w/o $\mathcal{P}^{T}$) sharply degrades the performance, which verifies the significance of textual prompting in learning targeted stance features for multi-modal stance detection. In addition, the removal of visual prompt tuning (w/o $\mathcal{P}^{V}$) leads to considerable performance degradation, which indicates that utilizing visual prompt tuning can make better learning of multi-modal targeted stance information and thus improves the performance of multi-modal stance detection.
Figure 3: Performance of using different pre-trained language models: Roberta Nguyen et al. (2020) (left) and KEBERT Kawintiranon and Singh (2021) (right). The reported results are the average results of all targets in one dataset on in-target multi-modal stance detection.
Figure 4: Performance of using different multi-modal fusion methods. The reported results are the average results of all targets in one dataset on in-target multi-modal stance detection.
6.4 Generalization of Targeted Multi-modal Prompt Tuning
Pre-trained Models
Previous experiments have demonstrated that the stance expression primarily resides in the textual modality. Therefore, to investigate the generalization of our TMPT when used with different pre-trained language models, we conduct experiments with two variants of our TMPT by using two other promising PLMs: RoBERTa Liu et al. (2019b) and KEBERT Kawintiranon and Singh (2021). The results are shown in Figure 3. Given the page constraints, we conduct experiments on Mtse, Mccq, and Mwtwt datasets. Note that our Targeted Multi-modal Prompting can directly work with the two PLMs and achieve better performance, showing the compatibility of our method with various pre-trained models.
Multi-modal Fusion
We choose cross-modal attention Wei et al. (2020) and GAT Velickovic et al. (2018) to analyze the performance of our TMPT with different multi-modal fusion methods. Given the page constraints, we conducted our experiments on Mtse, Mccq, and Mwtwt datasets. The results are shown in Figure 4. It can be seen that no matter which fusion method we use, the experimental performance is better than BERT+ViT. This indicates that our TMPT can directly work with various multi-modal fusion methods and lead to improved multi-modal stance detection performance.
Figure 5: Visualization of a typical example.
6.5 Visualization
To qualitatively investigate how our TMPT improves the performance of multi-modal stance detection, we visualize the attention values calculated by the targeted prompts and the vectors of the final layer of two encoders. The results are shown in Figure 5. It can be seen that the crucial textual tokens and the key visual patches regarding different targets are highly attended to and discriminated by our TMPT. This illustrates that the proposed TMPT can learn the important stance features for the specific target with the help of targeted prompt tuning, thus improving the learning ability of multi-modal stance detection.
7 Conclusion
This paper presents Mtse, Mccq, Mwtwt, Mruc and Mtwq, five new datasets for multi-modal stance detection. Based on the created datasets, we present in-target multi-modal stance detection and zero-shot multi-modal stance detection, aiming to advance and facilitate research in the field of multi-modal stance detection. In addition, we propose a simple yet effective targeted multi-modal prompting framework (TMPT) to deal with multi-modal stance detection, where the target information is explored to prompt the pre-trained models in learning multi-modal stance features. Extensive experiments on the new datasets show that our proposed TMPT achieves overall better performance than state-of-the-art baseline methods.
Limitations
Our method needs to devise a specific prompt for the given target, which needs to take time and artificial effort to analyze the target information in the real-world scenario for designing and selecting appropriate prompts. Furthermore, our proposed method does not integrate external target-specific knowledge to improve the learning of multi-modal stance information, such as the background knowledge of targets. Integrating external knowledge related to the target can improve the performance of stance detection. In addition, the current version of the data does not consider audio modality and video information, which is also an issue we need to explore in the future.
Ethics Statement
This work presents Mtse, Mccq, Mwtwt, Mruc and Mtwq, five new open-source datasets for the research community to study multi-modal stance detection. The Mtse, Mccq, and Mwtwt are the extension of Twitter Stance Election 2020 Kawintiranon and Singh (2021), COVID-CQ Mutlu et al. (2020), and Will-They-Won’t-They Conforti et al. (2020), which are three open-source textual stance detection datasets for academic research. We only collect the tweet text and image content needed for our research from Twitter following the privacy agreement of Twitter for academic usage, so there is no privacy issue. To annotate extended data, we recruited 8 experienced researchers who work on natural language processing or multi-modal learning. The detailed collect and annotate process has been illustrated in Section 3. Each researcher is paid \$6.5 per hour (above the average local payment of similar jobs). The entire annotation process lasted 5 months, and the average annotation time of the eight researchers was 430 hours. During the annotation process, samples that contain personally identifiable information will be discarded, and only the tweet IDs and human-annotated stance labels will be shared. Thus, our data set complies with Twitter’s information privacy policy. The annotators have no affiliation with any of the companies that are used as targets in the dataset, so there is no potential bias due to conflict of interest. We used the ChatGPT service from OpenAI for our writing. We followed their term and policies. Some examples in our paper may include a stance or tendency. It should be clarified that they are randomly sampled from the dataset for better studying the dataset and task, and do not represent any personal viewpoints.
Appendix A Keywords
In this section, we introduce the keywords to retrieve tweets.
A.1 Multi-modal Twitter Stance Election 2020
  • Since Twitter Stance Election 2020 Kawintiranon and Singh (2021) didn’t explicitly give the Keywords for collecting tweets, We used the keywords related to the election while with no clear preference: one of #vote, #Debates2020, #USElection2020, #PresidentialDebate2020, #2020Election, #votersuppression, #GetOuttheVote, #2020elections + mention of Trump or Biden.
  • Filter for posting time:
    Trump 01/01/2020 $\rightarrow$ 09/30/2020 Biden 01/01/2020 $\rightarrow$ 09/30/2020
A.2 Multi-modal COVID-CQ
  • We followed the keywords for collecting tweets from COVID-CQ Mutlu et al. (2020): one of hydroxychloroquine, chloroquine, HCQ.
  • Filter for posting time:
    Chloroquine 04/01/2020 $\rightarrow$ 04/30/2020
A.3 Multi-modal Will-They-Won’t-They
  • We followed the keywords for collecting tweets from Will-They-Won’t-They Conforti et al. (2020): one of merge, acquisition, agreement, acquire, takeover, buyout, integration + mention of a given company/acronym.
  • Filter for posting time:
    CVS_AET 02/15/2017 $\rightarrow$ 12/17/2018 CI_ESRX 05/27/2017 $\rightarrow$ 09/17/2018 ANTM_CI 04/01/2014 $\rightarrow$ 04/28/2017 AET_HUM 09/01/2014 $\rightarrow$ 01/23/2017 DIS_FOXA 07/09/2017 $\rightarrow$ 04/18/2018
A.4 Multi-modal Russo-Ukrainian Conflict
  • We used the keywords: Ukraine, Russia, Putin, Zelensky, Ukrainian, Russian
  • Filter for posting time:
    Russia 01/01/2022 $\rightarrow$ 06/30/2023 Ukraine 01/01/2022 $\rightarrow$ 06/30/2023
A.5 Multi-modal Taiwan Question
  • We used the keywords: Taiwan, Taiwan Crisis, Nancy Pelosi, Taiwan Strait
  • Filter for posting time:
    Mainland of China 01/01/2022 $\rightarrow$ 06/30/2023 Taiwan of China 01/01/2022 $\rightarrow$ 06/30/2023
Appendix B Annotation Guidelines
To ensure consistency with previous stance detection work, we follow the guidelines of Twitter Stance Election 2020 Kawintiranon and Singh (2021), COVID-CQ Mutlu et al. (2020), and Will-They-Won’t-They Conforti et al. (2020) to annotate the multi-modal stance of Mtse, Mccq and Mwtwt. For Mruc and Mtwq, the annotation guidelines are shown below:
B.1 Annotation Guidelines of MRUC and MTWQ
The annotation process consists of choosing one of three possible labels, given a tweet and an image. The three labels to choose from are Support, Oppose and Neutral.
Label 1: Support - If the tweet and image use direct or indirect expressions to support the target, or support objects which can represent the target (such as leaders, events).
Label 1: Oppose - If the tweet and image use direct or indirect expressions to oppose the target, or oppose objects which can represent the target (such as leaders, events).
Label 1: Neutral - If the tweet and image only mention the target without expressing a stance.
B.2 Details of Datasets Partition
As mentioned in Section 5.1, for each task/target, the dataset is divided into a training set, a development set, and a testing set with a ratio of 70%:10%:20%. In order to ensure that the partition is not affected by the data distribution bias. We performed 20 different random divisions for each task using the ratios above. We use the two baselines: BERT and BERT+ViT, to test every 20 groups of divisions. For each task, we take the division that can make the results of two baseline close to the median as the final partition.
Appendix C Prompt Tuning Analysis
C.1 Analysis of Textual Prompt Tuning
Frozen vs Tuned
For textual prompt tuning, we utilize a frozen paradigm to make full use of the semantic information learned by the pre-trained language model. To analyze the effectiveness of the frozen paradigm, we also tried a tuned prompt, the results are shown in Table 7. The results of the tuned paradigm are extremely poorer than the frozen one and fluctuate greatly. One possible reason is that, unlike the continuous picture patch in the visual transformer, the token in the textual modality is discrete single words, so it is hard to use a gradient-based method to fine-tune the prompts like a visual modality. Thus, in this paper, we choose a manually designed fixed prompt to obtain better performance.
Method Mtse Mccq Mwtwt Mruc Mtwq
Frozen 66.66 71.36 79.51 60.01 72.33
[…]
Table 7: Experimental results of fixed prompt and tuned prompt of textual modality. The reported results are the average results of all targets in one dataset on in-target multi-modal stance detection.
Hand-design textual prompts
To analyze the impact of the type of the hand-design textual prompts, we design several types of textual prompts and report the experiments in Table 8. Take the MTSE2020 dataset as an example, it can be seen that the experimental results of all types of prompts are superior to those without prompts (BERT+ViT). This demonstrates the significance of the targeted prompts in this task. Further, we can also see that there are considerable differences in performance between different types of prompts. Specifically, when using both target and stance to design prompts (③, ④, and ⑤), the model performs significantly better. Therefore, in our method, we choose the type of ⑤ to devise the textual prompts for multi-modal stance detection.
Method Textual Prompts Mtse
BERT+ViT - 63.28
Trump 63.32
Donald Trump 63.57
stance on Donald Trump 64.67
What is the stance on Donald Trump? 64.59
[…]
Table 8: Experimental results of using different textual prompts in MTSE dataset on target “Donald Trump”.
C.2 Analysis of Visual Prompt Tuning
Depth of prompt tuning
Following Jia et al. (2022), we conduct experiments with different depths of visual prompt tuning to analyze the impact of the depth of prompt tuning in our model. Given the page constraints, we conducted our experiments on MTSE, MCCQ, and MWTWT. The experimental results are shown in Table 9. We can see that both shallow and deep prompt tuning are good in our model, while the results of shallow prompt tuning are slightly better. Therefore, we adopt shallow prompt tuning in our method.
Method Mtse Mccq Mwtwt
Shallow (ours) 66.12 70.03 79.03
[…]
Table 9: Comparison results of using the shallow prompt tuning and deep prompt tuning. Shallow prompt tuning refers to only fine-tuning the prompt tokens in the Embedding layer of ViT, while deep prompt tuning refers to fine-tuning the prompt tokens in the Embedding layer of ViT and each layer in the Transformer. The reported results are the average results of all targets in one dataset on in-target multi-modal stance detection.
Number of visual prompt tokens
To analyze the impact of the number of visual prompt tokens, we set the value range of tokens as {3, 5, 7, 9} for comparative experiments. Given the page constraints, we conducted our experiments on MTSE, MCCQ, and MWTWT. The results are shown in Figure 6. Note that different values of tokens can have a certain impact on performance. When the value is 7, the overall performance of the model on all datasets is the best. Therefore, we set the number of tokens to 7 in our method.
Figure 6: Experimental results of using different numbers of visual prompting tuning tokens. The reported results are the average results of all targets in one dataset on in-target multi-modal stance detection.
Initialization of prompting tuning tokens
To analyze the impact of the initialization of prompting tuning tokens, similar to textual modality, we use target-related images to initialize prompt tokens. For example, on the Mtse dataset, we use the official portraits of two targets to initialize the corresponding visual prompting tuning tokens. The results are shown in Table 10. We can see that initializing prompting tuning tokens with images related to the target can achieve better results than random initialization.
Method Mtse
Initializing with related pictures (ours) 66.66
[…]
Table 10: Experimental results of using target-related images to initialize prompt tokens (ours) and random initialization. The reported results are the average results of all targets in one dataset on in-target multi-modal stance detection.
Appendix D Analysis of Our Datasets and Task
By comparing the results of BERT and ViT in Table 4 and 5, we ascertain that in the five multi-modal stance detection datasets, text modality is the dominant modality, comprising more stance information. Further, as described in Section 3.4 the annotators determine that nearly half of the samples contain stance information in their images, which maintains that the visual modality is important to facilitate textual modality in detecting stance, which can also be verified by the experimental results in Table 4 and 5 that BERT+ViT outperforms BERT and ViT and the visualization in Figure 5. In addition, from the results in Table 4 and 5, we can also see that even though our model performs unsatisfactorily on most datasets in all stance detection scenarios, which validates the challenge of our datasets and this task from another perspective.